OpenStudy (anonymous):

- 1/2 x + 3 = -x + 7 can someone explain how to solve this?

OpenStudy (luckyness):

Sure! :) But do you have an idea on what it is?

OpenStudy (luckyness):

\[\frac{ -1 }{ 2 }x+3=-x+7\] Okay so this is your equation. The first step would be to add x to both sides.

OpenStudy (anonymous):

@Luckyness It's either 8, -8, 2, or 1/2. I'm not really sure how to even begin solving this. :/

OpenStudy (jdoe0001):

hint: multiply both sides by 2

OpenStudy (luckyness):

No, you do both by x.

OpenStudy (luckyness):

Let me give you an attachment of the steps I am going off of.

OpenStudy (anonymous):

@Luckyness okay, thank you.

OpenStudy (luckyness):

Okay so these are the steps.

OpenStudy (luckyness):

So in the final step it says to solve this: \[2*(\frac{ 1 }{ 2 }x)=2*4\]

OpenStudy (luckyness):

So if you used a calculator and solved that, you would get x.

OpenStudy (luckyness):

Also, the answer was in one that you guessed. It is between 8 or 1/2. Your guess.

OpenStudy (anonymous):

@Luckyness thank you so much! & sorry for not responding, I was doing another question. I'm guessing that it is 8?

OpenStudy (luckyness):

Yes, it was 8. :)